1kg of ice at 20'c mixed with 2 kg water at 90'c.no los of energy to environment. What is the final temp of the mixture? Latent heat of ice=334.4kj/kg specific heat of water and ice are 4.18 and 2.09 kj/kgK resp

Grade:

1 Answers

Arun
6 years ago
 
Dear student
 
mi si \DeltaT + mi L + mi sw (T-0) = mw sw (90 -T)
1 * 2.09 *20 + 1 * 334.4 + 1 *4.18 * T =  2 * 4.18 * (90 -T)
 
T = 60 - 30 = 30 °c
 
Regards
Arun (askIITians forum expert)